OpenStudy (anonymous):

A ball is thrown horizontally from a 17m -high building with a speed of 7.0m/s . How far from the base of the building does the ball hit the ground? how do i approach this...?

OpenStudy (souvik):

the initial vertical velocity of the ball is zero.. let,it will take time t to hit the ground... using this...\[h _{y}=u _{y}+1/2g t^2\] here \(h _{y}\)=17 m \(u _{y}\)=0 take g=10 m/sec^2 so \[t^2=17*2/10\] the initial horizontal velocity of ball is 7 m/s and there is no acceleration in horizontal direction the distance from the base where the ball will hit the ground is \(v _{x}\)*t =\[7*\sqrt{34/10}\] ~13 m